Consider the vector space P3

5.1. Is span {1+x,x+x2,x2+x3,x3+1}=P3 Motivate your answer

5.2. Let D P3→ P2 be the differentiation operator D(a0+a1x+a2x2+a3x3)=a1+2a2x+3a3x2

(i) Find the matrix representation of D relative to the basis {1,x,x2,x3} using the coefficient ordering a0+a1x+a2x2+a3x3"\begin{bmatrix}
 a0 \\
 a1\\
 a2 \\
 a3
\end{bmatrix}"

(ii) Find the kernel and range of D


1
Expert's answer
2021-10-28T22:27:01-0400

PART 5.1


Relative to basis "(1,x,x\u00b2,x\u00b3)"

"(1+x)=(1,1,0,0),(x+x\u00b2)=(0,1,1,0)"

"(x\u00b2+x\u00b3)=(0,0,1,1),(x\u00b3+1)=(1,0,0,1)"

Matrix P="\begin{pmatrix}
 1& 0&0&1\\
 1& 1&0&0\\0&1&1&0\\0&0&1&1
\end{pmatrix}"


rref of P="\begin{pmatrix}
 1&0&0&1\\
 0&1&0&-1\\0&0&1&1\\0&0&0&0
\end{pmatrix}"


The rank of the matrix is 3, therefore it span P3


PART 5.2 (i)


"D(1)=0,D(x)=1,D(x\u00b2)=2x"


"D(x\u00b3)=3x\u00b2"

Relative to the basis

"0=(0,0,0,0), 1=(1,0,0,0)"

"2x=(0,2,0,0)"


"D=\begin{pmatrix}
 0 & 1&0&0 \\
 0&0&2&0\\0&0&0&3\\0&0&0&0
\end{pmatrix}"


Part 5.2 (ii)


rref "D=\begin{pmatrix}
 0&1&0&0\\
 0&0&1&0\\0&0&0&1\\0&0&0&0
\end{pmatrix}"


Range of D is 3


"\begin{bmatrix}
 0&1&0&0\\
 0&0&1&0\\0&0&0&1\\0&0&0&0
\end{bmatrix}\begin{bmatrix}
 x_1\\
 x_2\\x_3\\x_4
\end{bmatrix}=\begin{bmatrix}
 0\\
 0\\0\\0
\end{bmatrix}"


"x_2=0\quad x_3=0\quad x_4=0"

Taking "x_1=t\>where\>t\in\R"


Kernel "D=[{(t,0,0,0)\>:\>t\in\R}]"